# Iso-BAE Flows

### v1

@@ Abstract
We discover a family of local deformations that leave part of the spectrum intact for strongly interacting and exactly solvable quantum many-body systems. Since the deformation preserves the Bethe Ansatz equations (BAE), it is dubbed as the iso-BAE flow. Although all theories on the flow share the same BAE, the spectra are different. Part of the spectrum remains intact along the whole flow. Such states are protected by an emergent symmetry. The rest of the spectrum change continuously along the flow and are doubly degenerate for even length spin chains. For odd length chains, the deformed spectrum also comprises doubly degenerate pairs apart from the sector with magnon number $(L+1)/2$, where $L$ is the length of the spin chain. We discuss the iso-BAE flow for ${\rm XXX}_{1/2}$ model in detail and show that the iso-BAE flows exist for more general models including $q$-deformed XXZ as well as higher spin ${\rm XXX}_{s}$ spin chains.
@@ I Introduction
We will analyze the flow by two complementary approaches. In the first approach, we construct an operator $\mathcal{Q}_{\alpha}$ and show that it commutes with the Hamiltonian $[\mathbf{H}_{\alpha},\mathcal{Q}_{\alpha}]=0$ and is nilpotent (for even length) $\mathcal{Q}_{\alpha}^{2}=0$. It then follows that, once acting $\mathcal{Q}_{\alpha}$ on a given eigenstate $\mathcal{Q}_{\alpha}|\psi\rangle$ there are only two possibilities: either the result is vanishing or we obtain another eigenstate which is degenerate with $|\psi\rangle$, which correspond to type-I and type-II states, respectively. We then study more closely the eigenvalues by a careful scrutiny of BAE. To perform this analysis, we use the rational $Q$-system [10, 11, 12, 9, 13, 14], which gives complete physical spectrum. From the $Q$-system, we find that the results are also in accord with the classification of states mentioned before.
We first perform the analysis for the XXX1/2 type spin chain for simplicity. Then we will show that the construction can be generalized to the $q$-deformed and higher spin models. While each generalization has certain new features, the main characteristic, namely the undeformed BAE and the classification of type-I and type-II states remain the same as in the XXX1/2 case.
@@ II.1 Integrability
Naively (9) can be identified with the BAE of (2).  [^1] When $\alpha=-\beta=1$, the BAE for $(L-1)$ sites coincide with the BAE for $L$ site with $\beta=\alpha-1$. We explain the details in Sec. C of SM. However, one immediately sees a problem because the BAE (9) does not independ on the boundary parameter $\alpha$ while the Hamiltonian itself does ! In Sec. III, we will see that this is because (9) only describes part of the spectrum. Since the models (2) share the same BAE, we call the models with different $\alpha$ an iso-BAE flow.
@@ II.2 Type-I and type-II states
We will now show that for the model $\mathbf{H}_{\alpha}$, there exists an operator $\mathcal{Q}_{\alpha}$ commuting with Hamiltonian, i.e.
@@
$\mathcal{Q}_{\alpha}=\sum_{M=0}^{\lfloor L/2\rfloor]}\Pi_{L-M+1}\hat{\mathfrak{q}}_{\alpha}^{L-2M+1}\Pi_{M}\,.$ (11)
@@ II.3 Structure of the Hilbert space
For the undeformed model (corresponding to $\alpha=\infty$), which is SU(2) invariant, eigenstates form SU(2) multiplets. The highest weight states correspond to physical solutions of the BAE (9). Moreover, the operator $\hat{\mathfrak{q}}_{\alpha}$ is proportional to the global spin lower operator $\mathbf{S}^{-}$. Therefore, all the highest-weight states are type-I states, while all the descendants are type-II. In the $M$-magnon sector, there are $\mathcal{N}_{\text{I}}(L,M)=\begin{pmatrix}L\\ M\end{pmatrix}-\begin{pmatrix}L\\ M-1\end{pmatrix}$ physical solutions. Each solution correspond to a spin-$(\frac{L}{2}-M)$ representation with dimension $(L-2M+1)$.
@@
Figure 2: Lifting of the degeneracies for the spectrum for finite $\alpha$. The blue lines correspond to the energy eigenvalues for the type-I states, while the red lines are the type-II states. In this figure, the two type-I states are $M=2$ and $M=1$ for system size $L=6$ from bottom to up respectively. The type-II states are doubly degenerate in this case.
@@
The sector $M=(L+1)/2$ is special become in this sector $\mathcal{Q}_{\alpha}=\Pi_{(L+1)/2}$ and $\mathcal{Q}_{\alpha}|\psi_{(L+1)/2}\rangle^{\text{II}}=|\psi_{(L+1)/2}\rangle^{\text{II}}$.
@@ III BAE, $TQ$ and $QQ$ analysis
For $u=\pm\mathrm{i}(\alpha-\tfrac{1}{2})$, the right hand vanishes. If $Q(u)$ contains the roots $u=v_{M}=\pm\mathrm{i}(\alpha-\tfrac{1}{2})$, $TQ$-relation (18) can still be satisfied. Let us denote the rest of the Bethe roots by $\{v_{j}\}_{j=1}^{M-1}$, it follows that $v_{j}$ satisfy the following reduced BAE
@@ IV.1 XXZ model
The Hilbert spaces for even and odd lengths of the chain are decomposed precisely into the same way in (14) and (15), where the same doubly degeneracies for type-II states appear.
@@ IV.2 Higher spin $\mathrm{XXX}_{s}$ model
One important example of higher spin models is the spin-1 Takhtajan-Babujian model [22, 23], i.e. $\mathrm{XXX}_{1}$ model.
For $\mathrm{XXX}_{1}$ model, we can construct the following iso-BAE flow
@@ V Summary and discussions
In this work, we found an interesting deformation in quantum many-body system which keeps part of the spectra intact while deforming the rest. The undeformed spectra are protected by an emergent symmetry. Although easily stated, explicit constructions of such deformations have not been studied much. In the current work we rely on integrability, for obvious reasons: integrability offers us the necessary tools to find the complete spectrum. Careful analysis of the Bethe ansatz provides us clues to construct such flows. It would be very interesting to find examples with non-integrable models, but so far we are still lack of a clear guidance.
@@
Even if we restrict ourselves to integrable models, there are many questions to be answered. An immediate question is how general our construction is. The spin chain models we have considered in the current work are not the most general ones. More works are needed to see whether similar constructions exist for XXZ model when $q$ is a root of unity. There are some numerical evidences but the full structure is more subtle. Also, all the models we have considered so far preserve at least the U(1) symmetry, or magnon numbers. What about the models without U(1) symmetry such as the XYZ model [24, 25, 26] or models with un-parallel boundary magnetic fields [27, 28, 29, 30, 31, 32, 33]? Those questions remain elusive, and they might reveal the underlying algebraic structure of the iso-BAE flow.
@@
Going beyond spin chain models, it is natural to ask what about other types of models such as integrable field theories with boundaries. One possible way is considering proper continuum limit of the spin chain models [34, 35, 36, 37]. For example, in the case of ${\rm XXX}_{1/2}$ model, the field theory limit yields $c=1$ free compactified boson BCFT at self-dual radius $R=\sqrt{2}$ [15, 38]. The spectra of Neumann-Neumann and Dirichlet-Dirichlet boundary conditions coincide [39], which seems to match our expectation from iso-BAE flow. Following this route, it is highly likely that similar deformations exist for rather general integrable models.
@@
We would like to emphasis that although the spectrum of the type-I states are invariant under the flow, the eigenvectors are deformed. This implies that even for type-I states, other physical quantities such as correlation functions [40, 41] or entanglement measures [42, 43, 44] might vary with $\alpha$. Since the models along the iso-BAE flow is closely related to each other, it provides a new set-up to study quench dynamics [45, 46, 47] along the iso-BAE flow. Such process may be under better analytic control and shed new lights on the out-of-equilibrium dynamics of integrable spin chains.
Acknowledgment
@@
The work of Y.J. is partly supported by Startup Funding no. 3207022217A1 of Southeast University and by the NSF of China through Grant No. 12247103. Y.M.’s work is supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. We thank Jue Hou for previous collaborations on related topics. We are grateful to Philip Boyle Smith, Jie Gu, Hosho Katsura, Yifan Liu, Rafael Nepomechie, Bart Vlaar, Masahito Yamazaki and Xin Zhang for useful discussions. Y.M. would like to thank Lorentz Center for the hospitality.
Supplemental Material

### v2

# Spectrum-preserving Deformations of Integrable Spin Chains with Open Boundaries

@@ Abstract
We discover a family of local deformations that leave part of the spectrum intact for strongly interacting and exactly solvable quantum many-body systems. Since the deformation preserves the Bethe Ansatz equations (BAE), it is dubbed the iso-BAE flow. Although all theories on the flow share the same BAE, the spectra are different. Part of the spectrum remains intact along the whole flow. Such states are protected by an emergent symmetry. The remaining parts of the spectrum change continuously along the flow and are doubly degenerate for even length spin chains. For odd length chains, the deformed spectrum also comprises doubly degenerate pairs apart from the sector with magnon number $(L+1)/2$, where $L$ is the length of the spin chain. We discuss the iso-BAE flow for the ${\rm XXX}_{1/2}$ model in detail and show that the iso-BAE flows exist for more general models including $q$-deformed XXZ as well as higher spin ${\rm XXX}_{s}$ spin chains.
@@ I Introduction
We will analyze the flow by two complementary approaches. In the first approach, we construct an operator $\mathcal{Q}_{\alpha}$ and show that it commutes with the Hamiltonian $[\mathbf{H}_{\alpha},\mathcal{Q}_{\alpha}]=0$ and is nilpotent (for even length) $\mathcal{Q}_{\alpha}^{2}=0$. It then follows that, once acting $\mathcal{Q}_{\alpha}$ on a given eigenstate $\mathcal{Q}_{\alpha}|\psi\rangle$ there are only two possibilities: either the result is vanishing or we obtain another eigenstate which is degenerate with $|\psi\rangle$, which correspond to type-I and type-II states, respectively. We then study more closely the eigenvalues by a careful scrutiny of BAE. To perform this analysis, we use the rational $Q$-system [10, 11, 12, 9, 13, 14], which gives the complete physical spectrum. From the $Q$-system, we find that the results are also in accord with the classification of states mentioned before.
We first perform the analysis for the XXX1/2 type spin chain for simplicity. Then we will show that the construction can be generalized to the $q$-deformed and higher spin models. While each generalization has certain new features, the main characteristic, namely the undeformed BAE and the classification of type-I and type-II states, remains the same as in the XXX1/2 case.
@@ II.1 Integrability
Naively (9) can be identified with the BAE of (2).  [^1] When $\alpha=-\beta=1$, the BAE for $(L-1)$ sites coincide with the BAE for $L$ site with $\beta=\alpha-1$. We explain the details in Sec. C of SM. However, one immediately sees a problem because the BAE (9) does not depend on the boundary parameter $\alpha$ while the Hamiltonian itself does ! In Sec. III, we will see that this is because (9) only describes part of the spectrum. Since the models (2) share the same BAE, we call the models with different $\alpha$ an iso-BAE flow.
@@ II.2 Type-I and type-II states
We will now show that for the model $\mathbf{H}_{\alpha}$, there exists an operator $\mathcal{Q}_{\alpha}$ commuting with the Hamiltonian, i.e.
@@
$\mathcal{Q}_{\alpha}=\sum_{M=0}^{\lfloor L/2\rfloor}\Pi_{L-M+1}\hat{\mathfrak{q}}_{\alpha}^{L-2M+1}\Pi_{M}\,.$ (11)
@@ II.3 Structure of the Hilbert space
For the undeformed model (corresponding to $\alpha=\infty$), which is SU(2) invariant, eigenstates form SU(2) multiplets. The highest weight states correspond to physical solutions of the BAE (9). Moreover, the operator $\hat{\mathfrak{q}}_{\alpha}$ is proportional to the global spin lower operator $\mathbf{S}^{-}$. Therefore, all the highest-weight states are type-I states, while all the descendants are type-II. In the $M$-magnon sector, there are $\mathcal{N}_{\text{I}}(L,M)=\begin{pmatrix}L\\ M\end{pmatrix}-\begin{pmatrix}L\\ M-1\end{pmatrix}$ physical solutions. Each solution corresponds to a spin-$(\frac{L}{2}-M)$ representation with dimension $(L-2M+1)$.
@@
Figure 2: Lifting of the degeneracies for the spectrum for finite $\alpha$. The blue lines correspond to the energy eigenvalues for the type-I states, while the red lines are the type-II states. In this figure, the two type-I states are $M=2$ and $M=1$ for system size $L=6$ from bottom to top respectively. The type-II states are doubly degenerate in this case. Here we present eigenstates with different magnetization sectors, which are different from the ones in Fig. 1.
@@
The sector $M=(L+1)/2$ is special because in this sector $\mathcal{Q}_{\alpha}=\Pi_{(L+1)/2}$ and $\mathcal{Q}_{\alpha}|\psi_{(L+1)/2}\rangle^{\text{II}}=|\psi_{(L+1)/2}\rangle^{\text{II}}$.
@@ III BAE, $TQ$ and $QQ$ analysis
For $u=\pm\mathrm{i}(\alpha-\tfrac{1}{2})$, the right-hand side vanishes. If $Q(u)$ contains the roots $u=v_{M}=\pm\mathrm{i}(\alpha-\tfrac{1}{2})$, the $TQ$-relation (18) can still be satisfied. Let us denote the rest of the Bethe roots by $\{v_{j}\}_{j=1}^{M-1}$, it follows that $v_{j}$ satisfy the following reduced BAE
@@ IV.1 XXZ model
The Hilbert spaces for even and odd lengths of the chain are decomposed precisely into the same way in (14) and (15), where the same double degeneracies for type-II states appear.
@@ IV.2 Higher spin $\mathrm{XXX}_{s}$ model
One important example of higher spin models is the spin-1 Takhtajan-Babujian model [21, 22], i.e. the $\mathrm{XXX}_{1}$ model.
For the $\mathrm{XXX}_{1}$ model, we can construct the following iso-BAE flow
@@ V Summary and discussions
In this work, we found an interesting deformation in quantum many-body systems which keeps part of the spectra intact while deforming the rest. The undeformed spectra are protected by an emergent symmetry. Although easily stated, explicit constructions of such deformations have not been studied much. In the current work we rely on integrability, for obvious reasons: integrability offers us the necessary tools to find the complete spectrum. Careful analysis of the Bethe ansatz provides us clues to construct such flows. It would be very interesting to find examples with non-integrable models, but so far we still lack a clear guidance.
@@
Even if we restrict ourselves to integrable models, there are many questions to be answered. An immediate question is how general our construction is. The spin chain models we have considered in the current work are not the most general ones. More works are needed to see whether similar constructions exist for the XXZ model when $q$ is a root of unity. There are some numerical evidences but the full structure is more subtle. Moreover, the operators $\hat{\mathfrak{q}}_{\alpha}$ resemble the Yangian generators in the XXX case. The algebraic origin is yet to be elucidated. Also, all the models we have considered so far preserve at least the U(1) symmetry, or magnon numbers. What about the models without U(1) symmetry such as the XYZ model [23, 24, 25] or models with un-parallel boundary magnetic fields [26, 27, 28, 29, 30, 31, 32]? Those questions remain elusive, and they might reveal the underlying algebraic structure of the iso-BAE flow.
@@
Going beyond spin chain models, it is natural to ask about other types of models such as integrable field theories with boundaries. One possible way is considering the proper continuum limit of the spin chain models [33, 34, 35, 36]. For example, in the case of the ${\rm XXX}_{1/2}$ model, the field theory limit yields the $c=1$ free compactified boson BCFT at self-dual radius $R=\sqrt{2}$ [15, 37]. The spectra of Neumann-Neumann and Dirichlet-Dirichlet boundary conditions coincide [38], which seems to match our expectation from iso-BAE flow. Following this route, it is highly likely that similar deformations exist for rather general integrable models.
@@
We would like to emphasize that although the spectrum of the type-I states is invariant under the flow, the eigenvectors are deformed. This implies that even for type-I states, other physical quantities such as correlation functions [39, 40] or entanglement measures [41, 42, 43] might vary with $\alpha$. Since the models along the iso-BAE flow are closely related to each other, this provides a new set-up to study quench dynamics [44, 45, 46] along the iso-BAE flow. Such process may be under better analytic control and shed new light on the out-of-equilibrium dynamics of integrable spin chains.
Acknowledgements.
@@
The work of Y.J. is partly supported by Startup Funding No. 3207022217A1 of Southeast University and by the NSF of China through Grant No. 12247103. Y.M.’s work is supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. We thank Jue Hou for previous collaborations on related topics. We are grateful to Philip Boyle Smith, Jie Gu, Hosho Katsura, Yifan Liu, Rafael Nepomechie, Bart Vlaar, Masahito Yamazaki and Xin Zhang for useful discussions. Y.M. would like to thank Lorentz Center for hospitality.
Supplemental Material for
“Spectrum-preserving Deformations of Integrable Spin Chains with Open Boundaries”
Yunfeng Jiang and Yuan Miao